40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Nov 24 22:14:08 2012, Koji, Update, General, AP table cleaning up APTable121124.pdfAPTable121124.graffle.zip
    Reply  Tue Nov 27 03:26:00 2012, Jenne, Update, General, POP beam on POP55, POP camera 
    Reply  Tue Nov 27 17:45:42 2012, Steve, Update, General, spare optics of AP table moved to cabinet IMG_1819.JPG
    Reply  Wed Mar 28 16:44:27 2018, Steve, Update, General, AP table today AP_Table_20180328.png
Message ID: 7755     Entry time: Tue Nov 27 17:45:42 2012     In reply to: 7740
Author: Steve 
Type: Update 
Category: General 
Subject: spare optics of AP table moved to cabinet 

Spare optics from the AP table were moved to glass cabinet in the east arm. I'm not sure this is the right place. We'll see what everybody thinks.

There were two UNMARKED optics. Shame on you! No pencil marks on the optics either. These optics were shipped to the FBI for finger tip analysis.

 

 

Attachment 1: IMG_1819.JPG  1.810 MB  | Hide | Hide all
IMG_1819.JPG
ELOG V3.1.3-